OpenStudy (anonymous):

simplify r^2-9/(r+3)^2

OpenStudy (dumbcow):

r^2 - 9 is example of difference of squares --> a^2 - b^2 = (a+b)(a-b) think of 9 as 3^2

OpenStudy (anonymous):

sorry I don't understand what you are telling me

OpenStudy (anonymous):

so would the answer be r^2/2r

OpenStudy (dumbcow):

no what i was getting at is you can write (r^2 -9) as (r+3)(r-3)

OpenStudy (dumbcow):

then you should notice that something cancels ...

OpenStudy (anonymous):

ok so the r+3 cancles each other out.

OpenStudy (dumbcow):

yep...thus leaving the simplified answer :)

OpenStudy (anonymous):

r-3

OpenStudy (dumbcow):

over (r+3) remember it was squared, so only 1 cancelled

OpenStudy (anonymous):

ok thank you

OpenStudy (dumbcow):

(r+3)^2 = (r+3)(r+3) welcome